5/22-23: MZ release from PSL
5/20-24: Exchanging cryo valves.
5/27-6/2: Y arm mass lock. (not done yet)
5/29: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (specially BS)
6/3-6/8: Installation of TMS for ETMY
6/8: Engineering run
6/18 Move PR3
- 6/19 DRMI locked.
- 6/21: Stop cooler.
- 6/20: Stop cooler
- 7/1: Open chamber
- Long term:
4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- Please reduce the number or speed of 16kHz channels.
Report of work
- Type A, CRY: Because of slipping something ETMY did not come back, so we started warming up to open the chamber.
- Open day of ETMY chamber will be the next Friday.
- COM: Clipping was found in the PRMI cavity. It was about ~25%.
- ETMX outgas is increasing. We decided to stop cooler.
- Calibration of DRMI will be tested.